f(x)=1+2sin x+3cos ²x find it`s maximum value

Grade:Upto college level

1 Answers

Sher Mohammad IIT Delhi
10 years ago
f'(x)=2cosx-6cosxsinx=0
2cosx(1-3sinx)=0
cosx=0 or sin(x)=1 f(x)=1+2=3
sinx=1/3 imply cos^2(x)=1-1/9=8/9

f(x)=1+2/3+3*8/9=1+10/3=4.33
hence answer is 4.33.

Sher Mohammad
B.Tech, IIT Delhi.